Gas Laws. The volume of gas is inversely proportional to the pressure on the gas. If the volume is 6 cubic centimeters when the pressure on the gas is 8 kilograms per square centimeter, then what is the volume when the pressure is 12 kilograms per square centimeter?
Answer by stanbon(57361)

The volume of gas is inversely proportional to the pressure on the gas.
v = k/p
---------------
If the volume is 6 cubic centimeters when the pressure on the gas is 8 kilograms per square centimeter,
6 = k/8
k = 48
Equation:
v = 48/p
---------------
then what is the volume when the pressure is 12 kilograms per square centimeter?
v = 48/12
v = 4 cubic centimeters
=========================
Cheers,
Stan H.